# DriveGen3D: Boosting Feed-Forward Driving Scene Generation with Efficient Video Diffusion

### v1

@@ 4.1 Experimental Setup
For 3D scene reconstruction stage, we adopt novel view synthesis (NVS) to assess reconstruction quality, following the evaluation protocol established in Driving Forward [11]. Given sequential video frames from timestamps t1t-1 and $t+1$ as input, our model synthesizes the intermediate frame at timestamp $t$ for quantitative comparison against ground truth. We report Peak Signalto-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [54] in Table 1. To benchmark performance, we compare our DriveGen3D reconstruction model against the Driving Forward baseline (evaluate on ground truth images) across same validation frames and synthesized video outputs. This dual evaluation strategy disentangles reconstruction capability from generation artifacts, providing comprehensive insights into geometric and appearance recovery accuracy.
Implementation Details. We train the model with a resolution of $424\times 800$. We inference the model on NVIDIA H20 GPUs. When assessing the time cost of our proposed method, the baselines for the video generation model are MagicDriveDiT (17f) and MagicDriveDiT (233f). The 3D scene reconstruction stage is trained on NVIDIA H20 GPUs for 2 days.
@@
Table 2: Time cost of different attention blocks of MagicDriveDiT during inference.
@@
Table 3: Acceralerating the inference process of MagicDriveDiT. 17f and 233f denote the frames count of generated videos.
4.2 Experimental Results
4.2.1 Novel View Synthesis
Table 1 provides a comprehensive comparison of DriveGen3D against both optimization-based dynamic methods and feed-forward reconstruction methods. Notably, when using generated images instead of ground truth, DriveGen3D maintains competitive reconstruction quality with a PSNR of 22.84 and achieves the highest SSIM of 0.811, demonstrating strong temporal coherence and structure preservation in generated scenes. This suggests that despite operating on synthetic inputs, DriveGen3D produces reliable and structurally consistent 3D reconstructions, validating the effectiveness of its end-to-end pipeline.
4.2.2 Experimental Analysis and Ablation Study
Diffusion steps acceraleration. Firstly, we visualize the input differences and output differences in consecutive timesteps in Figure 5. It is observed that the default configurations of TeaCache, i.e. ALL, exhibits a U-shape for model output, with downward tend initially, nearly constant for the middle and upward until the end. The same is for the uncondition branch. The condition branch exhibits a slightly different phenomenon, with the start of model output much smaller. We also apply simple polynomial fitting to fit a relationship between model input and output and the use these cofficients to predict model output according to the input. As shown in the black lines in Figure 5, the quality of fitting is best for the condition branch, while ALL and uncondition do not fit well for the start. We attribute this to the different input and output relationships in the start of diffusion process. So we finally only apply TeaCache to the condition branch of MagicDriveDiT. As shown in Table 3, equiping MagicDriveDiT with TeaCache for condition branch has a speedup of nearly three times and two times for 17 frames and 233 frames. The perception metrics, mAP and mIoU only shows slight decrease.
Figure 4: Typical examples of tensors’ data distribution in different attention blocks of MagicDriveDiT.
(a) All
(b) Condition
(c) Uncondition
Figure 5: Visualization of input differences and output differences in consecutive timesteps. We plot all, condition and uncondition separetely.
@@ 4.2.2 Experimental Analysis and Ablation Study
Quantized DiT. Secondly, we profile the time cost of MagicDriveDiT. As shown in Table  2, cross-view attention is identified as the most computationally costly process in MagicDriveDiT. We further analyze different attention components in MagicDriveDiT and plot the distribution of $Q,K,V$, as depicted in Figure 4. Notably, $Q$ and $K$ exhibit a trend similar to that in Figure 4 of SageAttention2. An interesting phenomenon is observed: the numeric range of $V$ follows the order: spatial >> temporal >> cross-view. Specifically, the cross-view range is 10 times smaller than the spatial range and 5 times smaller than the temporal range. Given the principle of quantization, a smaller range is more conducive to quantization. Therefore, a promising solution is not only to apply SageAttention to cross-view attention but also substitute the high-precision quantization method for V with a lower-resolution one. Concretely, for $V$, the FP8, E4M3 data type can be replaced with FP8, E5M2 data type and INT8. We leave the latter for future attempt. In this paper, applying sageattention to the transformer blocks saves an additional 30 seconds (233 frames) of inference time after diffusion steps acceraleration while maintaining performance.
4.2.3 Visualization and Comparison
@@
Full pipeline results. In Figure 7, we show a typical output result of DriveGen3D. DriveGen3D can generate and reconstruct videos for more than 20S, 12FPS.
Figure 7: Visualization of multiview reconstruction video.
Comparison with raw videos. By default, our reconstruction stage takes the generated videos as input. We compare with DrivingForward to validate the difference of raw videos and generated videos. As evidenced in Table 1, DriveGen3D maintains fidelity in video generation that is essentially comparable to existing methodologies with GT images, exhibiting only a marginal degradation. Qualitative analysis is shown in Figure 1. In the front view image of reconstruction videos, the spatial location of generated bus is different from gt videos, we check the 3D Box layout and find the generation videos doesn’t match well. In the view of back left, the generation images is blurry, underscoring the needs of optimizing video generation model for reconstruction. Moreover, both GT videos and generation videos display drawbacks of flickering pole, pointing out the flaw of second stage reconstruction model.
5 Limitation and Future Work
While DriveGen3D demonstrates strong performance in efficient long-horizon video generation and rapid 3D scene reconstruction, several limitations remain. First, the reconstruction process relies on per-frame Gaussian representations without explicitly modeling object-level dynamics, which may limit fidelity in highly interactive or densely populated scenes. Second, although our feed-forward reconstruction achieves significant speedup, it may still suffer from minor temporal inconsistencies in low-texture regions or occluded areas.
In future work, we plan to explore semantic-aware 3D scene reconstruction that disentangles dynamic objects from static backgrounds, enabling object-level controllability and persistent identity tracking. Additionally, we aim to extend DriveGen3D toward full-scene temporal modeling by integrating recurrent or transformer-based consistency modules. Finally, enhancing reconstruction robustness under noisy video inputs or adverse weather conditions remains an important direction to improve real-world deployment readiness.
6 Conclusion
We present DriveGen3D, an efficient framework for synthesizing high-resolution, long-duration driving videos from textual descriptions and BEV layouts, and generating high-quality large-scale dynamic scenes. Our approach marks a significant advancement in world modeling by bypassing conventional voxel-based generation paradigms [21, 39] through a novel integration of longitudinal video generation and scene reconstruction modules. This architecture enables faithful reproduction of real-world driving scenarios and thus paves the way for novel applications in autonomous vehicle simulation and dynamic world modeling.

### v2

@@ IV-A Experimental Setup
Figure 3: Typical examples of the data distribution of tensors in different attention blocks of MagicDriveDiT.
(a) All
(b) Condition
(c) Uncondition
Figure 4: Visualization of input differences and output differences in consecutive timesteps. We plot all, condition and uncondition separately.
@@
For 3D scene reconstruction stage, we adopt the view synthesis task to assess reconstruction quality, following the evaluation protocol established in Driving Forward [3]. Given sequential video frames from timestamps $t-\Delta$ and $t+\Delta$ (In our implementation, $\Delta=2$) as input, our model synthesizes the intermediate frame at timestamp $t$ for quantitative comparison against ground truth. We report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [32] in Table I.
TABLE II: Time cost of different attention blocks of MagicDriveDiT during inference.
@@
TABLE III: Acceralerating the inference process of MagicDriveDiT. 17f and 233f denote the frames count of generated videos.
@@
To benchmark performance, we compare our DriveGen3D reconstruction model against the Driving Forward baseline (evaluate on ground truth images) across same validation frames and synthesized video outputs. This dual evaluation strategy disentangles reconstruction capability from generation artifacts, providing comprehensive insights into geometric and appearance recovery accuracy.
Implementation Details. We train the model with a resolution of $800\times 424$. Inference is performed on NVIDIA H20 GPUs. When assessing the time cost of our proposed method, the baselines for the video generation model are MagicDriveDiT [1](17f) and MagicDriveDiT [1](233f). The 3D scene reconstruction stage is trained on a single NVIDIA H20 GPU for 2 days.
IV-B Experimental Results
IV-B1 Novel View Synthesis
Table I provides a comprehensive comparison of DriveGen3D against both optimization-based dynamic methods and feed-forward reconstruction methods. Notably, when using generated images instead of ground truth, DriveGen3D maintains competitive reconstruction quality with a PSNR of 22.84 and achieves the highest SSIM of 0.811, demonstrating strong temporal coherence and structure preservation in generated scenes. This suggests that despite operating on synthetic inputs, DriveGen3D produces reliable and structurally consistent 3D reconstructions, validating the effectiveness of its end-to-end pipeline.
IV-B2 Experimental Analysis and Ablation Study
We construct the following experiments to demonstrate the effectiveness of FastDrive-DiT.
Diffusion steps acceraleration. Firstly, we visualize the input differences and output differences in consecutive timesteps in Figure 4. It is observed that the default configurations of TeaCache [9], i.e. all, exhibits a U-shape for model output, with downward tend initially, nearly constant for the middle and upward until the end. The same is for the uncondition branch. The condition branch exhibits a slightly different phenomenon, with the start of model output much smaller. We also apply simple polynomial fitting to fit a relationship between model input and output and the use these cofficients to predict model output according to the input.
@@ IV-B2 Experimental Analysis and Ablation Study
Figure 6: Comparison of GT videos and generated video for reconstruction.
As shown in the black lines in Figure 4, the quality of fitting is best for the condition branch, while all and uncondition do not fit well for the start. We attribute this to the different input and output relationships in the start of diffusion process. So we finally only apply TeaCache to the condition branch of MagicDriveDiT. As shown in Table III, equiping MagicDriveDiT with TeaCache for condition branch has a speedup of nearly $3\times$ and $2\times$ for 17 frames and 233 frames. The perception metrics, mAP and mIoU only shows slight decrease.
Quantized DiT. Secondly, we profile the time cost of MagicDriveDiT. As shown in Table  II, cross-view attention is identified as the most computationally costly process in MagicDriveDiT. We further analyze different attention components in MagicDriveDiT and plot the distribution of $Q,K,V$, as depicted in Figure 3. Notably, $Q$ and $K$ exhibit a trend similar to that in Figure 4 of SageAttention2 [11]. An interesting phenomenon is observed: the numeric range of $V$ follows the order: spatial >> temporal >> cross-view. Specifically, the cross-view range is 10 times smaller than the spatial range and 5 times smaller than the temporal range. Given the principle of quantization, a smaller range is more conducive to quantization. Therefore, a promising solution is not only to apply SageAttention [10] to cross-view attention but also substitute the high-precision quantization method for V with a lower-resolution one. Concretely, for $V$, the FP8, E4M3 data type can be replaced with FP8, E5M2 data type and INT8. We leave the latter for future attempt. In this paper, applying sageattention to the transformer blocks saves an additional 30 seconds (233 frames) of inference time after diffusion steps acceraleration while maintaining performance.
IV-B3 Visualization and Comparison
We provide visual results of DriveGen3D alongside ablation experiments.
@@ IV-B3 Visualization and Comparison
Full pipeline results. In Figure 2, we show a typical output result of our framework. DriveGen3D can generate and reconstruct street scenes for more than 20s, 12FPS.
Comparison with raw videos. By default, our reconstruction stage takes the generated videos as input. We compare with DrivingForward to validate the difference of raw videos and generated videos. As evidenced in Table I, DriveGen3D maintains fidelity in video generation that is essentially comparable to existing methodologies with GT images, exhibiting only a marginal degradation. Qualitative analysis is shown in Figure 6. In the front view image of reconstruction videos, the spatial location of generated bus is different from gt videos, we check the 3D Box layout and find the generation videos does not match well. In the view of back left, the generation images is blurry, underscoring the needs of optimizing video generation model for reconstruction. Moreover, both GT videos and generation videos display drawbacks of flickering pole, pointing out the flaw of second stage reconstruction model.
V Conclusion
We present DriveGen3D, an efficient framework for synthesizing long-duration driving videos from textual descriptions and BEV layouts, and generating high-quality large-scale dynamic scenes. Our approach marks a significant advancement in world modeling by bypassing conventional voxel-based generation paradigms [21] through a novel integration of longitudinal video generation and scene reconstruction modules. This architecture enables faithful reproduction of real-world driving scenarios and thus paves the way for novel applications in autonomous vehicle simulation and dynamic world modeling.